OpenStudy (anonymous):

what is 5.6 x 10^-3 yd/year in feet per day

OpenStudy (jdoe0001):

\(\bf \cfrac{5.6 \times 10^{-3}\cancel{ yard }}{\cancel{ year }}\times \cfrac{3feet}{\cancel{ yard }}\times \cfrac{\cancel{ year }}{360days}\implies \square ?\)